OpenStudy (anonymous):

The position of an object as a function of time is given x=bt^2 -ct , where b=2.0 m/s^2 and c=6.7 m/s, and x and t are in SI units. What is the instantaneous velocity of the object when t=2.2?

OpenStudy (anonymous):

So x=2t^2-6.7t. When t=2.2, x=2(2.2)^2-6.7*2.2 x=-5.06...

OpenStudy (anonymous):

Oh wait

OpenStudy (anonymous):

Sorry

OpenStudy (anonymous):

You need the derivative

OpenStudy (anonymous):

x=2t^2-6.7t. The derivative is 4t-6.7 So the velocity is 2.1 m/s

OpenStudy (anonymous):

This is Calc right?

OpenStudy (anonymous):

this right answer

OpenStudy (anonymous):

Calc-Based Physic.)